OpenStudy (anonymous):

What are the real or imaginary solutions of each polynomial equation? 125x+343=0

OpenStudy (whpalmer4):

I bet that is \[125x^3+343=0\]

OpenStudy (whpalmer4):

which is a sum of cubes: \((5x)^3 + (7)^3\) Factor sum of cubes: \[a^3+b^3= (a+b)(a^2-ab+b^2)\]

OpenStudy (whpalmer4):

\[(5x+7)((5x)^2-(5x)(7) + (7)^2) = 0\]\[(5x+7)(25x^2-35x+49) = 0\] \[5x+7=0\]\[x=-\frac{7}{5}\] \[25x^2-35x+49=0\]Using quadratic formula:\[x = \frac{-(-35)\pm\sqrt{(-35)^2-4(25)(49)}}{2*25}\]\[x = \frac{7}{10}(1\pm i\sqrt{3})\]